2019 Account Cost Retail Inv, Jan 1 $ 4,000 9,000 Purchases 11,000 19,000 Freight In 2,000...

2019

Account

Cost

Retail

Inv, Jan 1

$ 4,000

9,000

Purchases

11,000

19,000

Freight In

2,000

Markups

1,000

Markup Cancelations

300

Markdowns

500

Markdown Cancelations

100

  

   Net Sales

14,000

   Employee Discounts

1,500

   Normal Shortage

500

  1. Using the Conventional Method, what is the Cost-to-Retail Ratio for 2019?
  2. Using the Conventional Method, what is the ending inventory valued at Cost for 2019?
  3. Using the LIFO Retail Method, what is the Cost-to-Retail Ratio for 2019?

Show how you solved it.
Answer #1
Conventional Method
Cost price Retail price
Opening Inventory        4,000 9,000
Add Purchases      11,000 19,000
Freight in        2,000
Net markups 700
(Markups - Markup cancellations) = $1000-$300 = $700
Total      17,000      28,700
A B
a. Cost to retail ratio = A/B 59.23%
Less: Net Markdowns = C -400
Total at retail = (B-C)      28,300
Less Sales -14000
Employee discounts -1500
Normal shortage -500
Closing Inventory at retail      12,300
b. Closing inventory at cost = Closing inventory at retail * cost to retail ratio        7,286
LIFO Retail Method
Cost price Retail price
Opening Inventory not to be considered not to be considered
Add Purchases        11,000            19,000
Freight in          2,000
Net markups 700
Less Net Markdowns = (Markdowns - Markdown cancellation)=$500-$100=$400 -400
Total        13,000            19,300
A B
c. Cost to retail ratio = A/B 67.36%